OpenStudy (anonymous):

Rearrange X^3-x^2-2x+1=0 in the form 1/x+x^2=ax+b where a and b are integers.

OpenStudy (anonymous):

X^3-x^2-2x+1=0 1+x^3=x^2+2x x(1/x+x^2)=x(x+2) (1/x+x^2)=x+2

OpenStudy (anonymous):

Can you give step by step how to do it?

OpenStudy (anonymous):

X^3-x^2-2x+1=0 Can u take x common from here

OpenStudy (anonymous):

Congratulations @sauravshakya and @ghazi for Golden..

OpenStudy (anonymous):

Thanx @waterineyes

OpenStudy (anonymous):

What do you mean by x common?

OpenStudy (lgbasallote):

congratulations @lgbasallote for beating @saifoo.khan finally

OpenStudy (ghazi):

@waterineyes thanks....

OpenStudy (anonymous):

X^3-x^2-2x+1=0 x(x^2-x-2+1/x)=0

Parth (parthkohli):

@lgbasallote that's epic, really!

OpenStudy (anonymous):

I just took x common from X^3-x^2-2x+1=0

OpenStudy (anonymous):

Welcome to both..

OpenStudy (ghazi):

\[x^3-x^2-2x+1=0....x(x^2-x-2+\frac{ 1 }{ x })\]

OpenStudy (ghazi):

@waterineyes got to know about you...

OpenStudy (anonymous):

@Ausam ????

OpenStudy (anonymous):

What @ghazi ??

OpenStudy (ghazi):

you're a girl.. :)

OpenStudy (anonymous):

Sorry. Press by mistake

OpenStudy (anonymous):

Having any doubt ??

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

Ha ha ha ha...

OpenStudy (ghazi):

before lookin at your profile...i was in doubt..but first solve this doubt

OpenStudy (anonymous):

I still don't get how to do it clearly

OpenStudy (anonymous):

I am asking to @ghazi Anyways what is your doubt??? @asdfasdfasdfasdfasdf1

OpenStudy (ghazi):

@Ausam take out x from equation

OpenStudy (anonymous):

X^3-x^2-2x+1=0 x(x^2-x-2+1/x)=0 x^2-x-2+1/x=0

OpenStudy (anonymous):

First clear the doubt of @asdfasdfasdfasdfasdf1 then we will talk..

OpenStudy (ghazi):

okay

OpenStudy (anonymous):

Now, arrange the left hand side as the question............ ANd took the rest to right hand side

OpenStudy (ghazi):

@Ausam divide whole equation by X

OpenStudy (anonymous):

Thanks. Thatd be great

OpenStudy (anonymous):

Oh really sorry gotta go now I will clear your doubt next time I will come online.. Sorry for that.. I am in college computer lab and now it is time to close the lab.. So you have to wait till tomorrow.. @ghazi

OpenStudy (ghazi):

it is cool

OpenStudy (anonymous):

So I factorize using x nd then arrange it like the equation given?

OpenStudy (ghazi):

\[\frac{ x^3-x^2-2x+1 }{ x }=\frac{ x^3 }{ x }-\frac{ x^2 }{ x }-\frac{ 2x }{ x }+\frac{ x }{ x }\] @Ausam clear?

OpenStudy (anonymous):

Yeah. Thanks

OpenStudy (ghazi):

\[x^2-x-2+\frac{ 1 }{ x }\] is it clear now?

OpenStudy (ghazi):

now you to do the separation...that's it

OpenStudy (anonymous):

Thanks again

OpenStudy (ghazi):

:)